OpenStudy (anonymous):

Graph the Rational Function f(X) = -X^2-7X-4/X=3. Please show graphing points

OpenStudy (anonymous):

x+3 not x=3 sorry

OpenStudy (anonymous):

\[f(x) = -x^{2}-7x-\frac{4}{x}-3\] Is that the equation? Or is it this? \[f(x) = -x^{2}-7x-\frac{4}{x-3}\] Or something else?

OpenStudy (anonymous):

\[-x^2-7x-4\div x+3\]

OpenStudy (anonymous):

so \[-x^2 - 7x -4\] ---------- \[x+3\]

OpenStudy (anonymous):

FYI, click equation below and use frac{}{}, example: frac{-x^{2}-7x-4}{x+3} turns into: \[\frac{-x^{2}-7x-4}{x+3}\] This is one of those times where there's not much else to do but to put it into a graphing calculator: http://www.wolframalpha.com/input/?i=%28%E2%88%92x%5E2%E2%88%927x%E2%88%924%29%2F%28x%2B3+%29

OpenStudy (anonymous):

well you could try estimating the positions of the curve to get a general shape of the actual graph using an interval chart.